A feed store sells two types of cattle feed. One sells for $180 and the other sells for $210 per ton. If the store wants to sell 60 tons of a mix that can be sold for $200 per ton, how much of each feed should be mixed?
Answer by jorel555(1290)

Let n be the amount of more expensive feed. Then the cheaper feed would be 60-n. So:
210n+180(60-n)=200(60)
30n+10800=12000
30n=1200
n=40
The store needs 40 tons of $210 feed, and 20 tons of $180 feed for its' mix. ☺☺☺☺